How fast will a 4 meter long car need to go in order to fit inside of a 3 meter long garage? (Give your answers in terms of c, the speed of light)


1
Expert's answer
2022-01-16T13:08:59-0500

Apply the equation for length contraction:


"L=L_0\sqrt{1-(v^2/c^2)}→v=c\dfrac{L}{L_0},\\\space\\
v_c=3/4=0.75c."